OpenStudy (anonymous):

Solve the equation: 4x2=9

OpenStudy (anonymous):

\[4x^2=9\] \[x^2=\frac{9}{4}\] \[x=\pm\frac{3}{2}\]

OpenStudy (anonymous):

Whoa Satellite73, that speed!